# Anchors Aweigh! Sail for Optimal Unified Multi-Modal Representations

### v2

@@ C.2 Experiments with real-world datasets
We utilize Low-Rank Adaptation (Hu et al., 2022) for training CentroBind and FABind , enhancing training efficiency and achieving impressive results with fewer iterations. For parameter settings, we set a learning rate of $0.001$, the AdamW optimizer (Loshchilov & Hutter, 2019) with a batch size of $16$, and a temperature of $0.3$ for InfoNCE. Training CentroBind requires augmentation. We augment video frames with various transformations, including random perspective shifts, random flips and rotation, color jitter, Gaussian blur, and auto-contrast adjustment. For the audio modality, we apply a low-pass filter, speed changes, echo effect, room impulse response convolution, and background noise. For the text modality, we generate paraphrased sentences using the Phi-3 language model served using Ollama  [^5] https://ollama.com/library/phi3.
@@ C.3 Experiments with pre-trained backbone
Figure 11: Top-1 retrieval accuracy of ImageBind and CentroBind on DreamBooth datset.
In addition to experiments on synthetic and real-world datasets, we evaluate our method using a pre-trained ImageBind backbone for image and text modalities. We utilize the DreamBooth dataset (Ruiz et al., 2023), which comprises 30 subjects across 15 different classes of image-text, to fine-tune the pre-trained encoders. The evaluation is conducted through a retrieval task, effectively measuring the quality of the learned unified embedding space.
Specifically, we begin by loading the pre-trained ImageBind model (Girdhar et al., 2023) and fine-tuning its text and image encoders on the DreamBooth training dataset using both the ImageBind and CentroBind methods. For fine-tuning, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2022) with a rank of 4, resulting in 5.4 million trainable parameters out of a total of 1.2 billion parameters. After fine-tuning, we measure the top-1 retrieval accuracy on the test dataset. The optimization process uses the AdamW optimizer with the following hyperparameters: a batch size of 16, a learning rate of $5.0\times 10^{-6}$, a maximum of 500 epochs, a weight decay of $10^{-4}$, and a temperature parameter set to 0.07. In this experiment, we omit the text augmentation step because the text samples in the DreamBooth dataset are too short. Instead, we create a centroid-based dynamic anchor using the embeddings of augmented images and original text.
The results presented in Figure 11 demonstrate that CentroBind outperforms the ImageBind method, even when leveraging a powerful pre-trained backbone and the image modality. We anticipate that this performance gap would widen if the backbone were less powerful or if the image modality were excluded. These findings corroborate our analysis of the dynamic anchor binding method, highlighting its effectiveness in enhancing multi-modal representations.

### v3

@@ C.2 Experiments on MUStARD
We utilize Low-Rank Adaptation (Hu et al., 2022) for training CentroBind and FABind, enhancing training efficiency and achieving impressive results with fewer iterations. For the backbones in FABind and CentroBind, we use the pretrained VideoMAE (Tong et al., 2022) for video data, the pretrained WaveLM (Chen et al., 2022) for audio data, and the pretrained BERT (Devlin et al., 2019) for text data. For parameter settings, we set a learning rate of $0.001$, the AdamW optimizer (Loshchilov & Hutter, 2019) with a batch size of $16$, and a temperature of $0.3$ for InfoNCE. Training CentroBind requires augmentation. We augment video frames with various transformations, including random perspective shifts, random flips and rotation, color jitter, Gaussian blur, and auto-contrast adjustment. For the audio modality, we apply a low-pass filter, speed changes, echo effect, room impulse response convolution, and background noise. For the text modality, we generate paraphrased sentences using the Phi-3 language model served using Ollama  [^4] https://ollama.com/library/phi3.
@@ C.3 Experiments with pre-trained backbone
(a) Image-text pair: DreamBooth (Ruiz et al., 2023)
(b) Image-audio pair: AVE (Tian et al., 2018)
(c) Image-audio pari: AudioSet (Gemmeke et al., 2017)
(d) Image-audio pair: UR-FUNNY (Hasan et al., 2019)
(e) Image-text pair: UR-FUNNY (Hasan et al., 2019)
(f) Image-audio-text pair: Dream+AVE+AudioSet
Figure 11: We evaluate cross-modal retrieval using the pretrained ImageBind backbone, fine-tuned on each dataset with both ImageBind and CentroBind. Apart from the image–text pair in the UR-FUNNY dataset—where performance is identical—CentroBind surpasses ImageBind in every other setting.
So far, the superiority of CentroBind is verified for controllable synthetic and MUStARD datasets. Although such experiments and theoretical analysis show the effectiveness of CentroBind, we further evaluate and compare the performance of CentroBind and FABind to validate in additional scenarios. In particular, we consider the following cases:
  1. 1.
    Bi-modal scenario $\rightarrow$ We compare performance in bi-modal datasets, DreamBooth (Ruiz et al., 2023), AVE (Tian et al., 2018), AudioSet (Gemmeke et al., 2017), UR-FUNNY (Hasan et al., 2019).
  2. 2.
    When a large-scale powerful backbone is available $\rightarrow$ We compare performance given the pre-trained ImageBind backbone.
  3. 3.
    When strong anchor modality is available $\rightarrow$ We compare performance with the image modality as anchor.
We conduct fine-tuning the ImageBind and CentroBind on DreamBooth (Ruiz et al., 2023), AVE (Tian et al., 2018), AudioSet (Gemmeke et al., 2017), UR-FUNNY (Hasan et al., 2019) datasets. For datasets containing video modality instead of image, we extract the middle frame from the video and use the middle frame as image sample. Audio is also extracted from video if audio modality does not exist in original dataset. The evaluation is conducted through a retrieval task, effectively measuring the quality of the learned unified embedding space.
After fine-tuning, we measure the top-1 and top-5 retrieval accuracy on the test dataset. The optimization process uses the AdamW optimizer with the following hyperparameters: a batch size of 16, a learning rate of $5.0\times 10^{-6}$, a weight decay of $10^{-4}$, and a temperature parameter set to $0.07$. We fine-tune the models until their validation accuracies converge to certain value. In this experiment, we omit the text augmentation. Instead, we create a centroid-based dynamic anchor using the embeddings of augmented images and original text.
More specifically, we begin by loading the pre-trained ImageBind model (Girdhar et al., 2023) and fine-tune its encoders. For fine-tuning, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2022) with a rank of 4, resulting in 5.1-5.4 million trainable parameters (depending on chosen modalities) out of a total of 1.2 billion parameters. For the experiment on AudioSet, we fine-tune the models on the Balanced train subset of AudioSet, provided in official site. For datasets provided without train-test split, we randomly split the dataset with $0.8:0.2$ ratio before finetuning. The results in Figure 11 and Table 3 show that CentroBind outperforms ImageBind even with a strong pretrained backbone and the image modality. Moreover, CentroBind achieves additional gains in bimodal settings, suggesting it mitigates FABind ’s inherent limitations across most scenarios. We anticipate that the performance gap would widen if the backbone were weaker, if the image modality were excluded, or if additional modalities were introduced. These findings corroborate our analysis of dynamic anchor binding, highlighting its effectiveness in enhancing multimodal representations.